OpenStudy (anonymous):

help intergrate from x=0 to x=1, 1 devide by [(x+1)powerd by 2][((x powered by 2)+ 1]

OpenStudy (experimentx):

\[\int\limits_{}^{}1/((x+1)^2(x^2+1) dx\]

OpenStudy (anonymous):

that's it, how did you write it simply like that ?

OpenStudy (experimentx):

click on Equation below your textbox .. left corner

OpenStudy (anonymous):

so do you have the solution ?

OpenStudy (anonymous):

partial fractions big pain in the neck

OpenStudy (anonymous):

can you do it ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but are you able to work it out ?

OpenStudy (anonymous):

yes, idea is to write \[\frac{1}{(x+1)^2(x^2+1)}=\frac{a}{x+1}+\frac{b}{(x+1)^2}+\frac{cx+d}{x^2+1}\]

OpenStudy (anonymous):

how is that simplified ?

OpenStudy (anonymous):

then solve for the coefficients a, b, c and d. it is a big pain in the neck. will get a system of linear equations to solve i would cheat and look here http://www.wolframalpha.com/input/?i=1%2F%28%28x%2B1%29%5E2%28x%5E2%2B1%29%29 where is says "partial fraction expansion" if you want the method click on "show steps"

OpenStudy (anonymous):

once you see it you will realize why i said "no"

OpenStudy (anonymous):

how about \[\int\limits_{?}^{?}1\div(5+3cosx) dx\]

OpenStudy (anonymous):

help